What is the value of v in the formula v=23t when t=1 1/4? 2 1/2? 3 1/4? 5 1/4? 5 1/2?
Answer by stanbon(75887)

What is the value of v in the formula v=23t when t=1 1/4?
-----
v = 23(1 1/4) = 23(5/4) = 28.75
-----------
Cheers,
Stan H.
-------